OpenStudy (anonymous):

x2 + 7x + 12 find the solution. a) x = -4, x = 3 b) x = -3, x = 4 c) x = 3, x = 4 d) x = -4, x = -3

OpenStudy (anonymous):

d) x = -4, x = -3

OpenStudy (anonymous):

2x+7x=-12 9x=-12 9x/9=-12/9 x=-4/3

OpenStudy (radar):

You have presented an expression not an equation. Therefore, it can not be solved. Is there more to the problem? Like does the above expression equal something. If it equals zero, then it would factor quite nicely into (x+4)(x+3)=0 giving solutions of x=-4, x=-3.